OpenStudy (anonymous):

If an object is dropped from a height of 144 feet, the function h(t)=-16t2t+144 gives the height of the object after t seconds. When will the object hit the ground? A. 9 s B. 6 s C. 1.5 s D. 3 s

OpenStudy (anonymous):

what it the position fundtion what is -162t2 ? do you mean \[-16t ^{2}+t\]

OpenStudy (anonymous):

ok set it equal to zero and solve for t.

OpenStudy (anonymous):

check it @Mr.Me

OpenStudy (anonymous):

-16t^2+144=0 16t^2=-144 t^2=9 ...

OpenStudy (anonymous):

so t=+- 3 but -3 doesn't make sense so t=3s is good answer

OpenStudy (anonymous):

so is it A or D

OpenStudy (anonymous):

d because t^2=9 and t=3

OpenStudy (anonymous):

okay thanks